OpenStudy (anonymous):

Accounting II question 36,000 - 24, 000 - 16,800= the ratio is 2:1 what would the solution be?

OpenStudy (anonymous):

Please put up the whole question. There is not enough here.

OpenStudy (gw2011):

Is this what you are looking for: 36,000-24,000-16,800=-4800 If the ratio is 2:1, then I assume you are looking for the following: -4800/x=2/1 2x=-4800 x=-2400 This will give you a ratio of 2:1

OpenStudy (anonymous):

Problem Alex Hutchins and Catlin Jenkins formed a partnership,dividing income as follow: 1. annual salary allowance of hutchins of 24,000 2 Interest of 12% on each partners capital balance of january 1 3. any remaining net income divided to hutchins and jenkins, 2:1 hutchins and jenkins had 60,000 and 80,000 respectively in their jan 1 capital balances. net income for the year was 36,000. How much net income should be distributed to Hutchins?

OpenStudy (anonymous):

is the 36,000 net income before salary and interest or after?

OpenStudy (anonymous):

if b4, then 36,000 - (24,000 + (12% of 60,000)) = 36000 - 31,200 = 4,800 net to hutchins will then be 2/3 x 4800 = 3200 if after, then net to hutchins = 2/3 x 36,000 = 24,000